Pete's Power Pizzas are microwaveable low-fat pastry shells filled with natural and nutritious ingredients, designd for the health-conscious person on the go. Pete's has determined that the demand function for the "Original" Power Pizza is: Quantity = 245.11 - 27.97 x Price Find Maximum Reservation Price for the "Original" Power Pizza.
Answer #1

Given :Demand function is ,

Quantity =245.11-27.97*Price

Maximum reservation price (MRP) is the price at which the quantity demanded =0

i.e 245.11-27.97*Price =0

Price=245.11/27.97 =8.76332

Maximum reservation price =8.76 (rounded to two decimal places)